California​ Gardens, Inc.,​ prewashes, shreds, and distributes a variety of salad mixes in​ 2-pound bags. Doug​ Voss, Operations​ VP, is considering a new​ Hi-Speed shredder to replace the old​ machine, referred to in the shop as​ "Clunker." Hi-Speed will have a fixed cost of ​$83 comma 000 per month and a variable cost of ​$1.25 per bag. Clunker has a fixed cost of only ​$43 comma 000 per​ month, but a variable cost of ​$1.75. Selling price is ​$2.50 per bag. ​a) What is the crossover point in units​ (point of​ indifference) for the​ processes? The crossover point is nothing units. ​(Round your response to the nearest whole​ number.)
Answer #1

Let the point of indifference be X, then

Total Expense of High Speed = Total Expense of Clunker

83000 + 1.25*X = 43000 + 1.75*X

0.5*X = 40000

X = 80000 Units